OpenStudy (anonymous):

What is the product of three over four times eight over fifteen in simplest form? one over twenty two over five thirty-two over forty-five twenty-four over sixty

OpenStudy (anonymous):

@tester

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I mean 2/5

OpenStudy (anonymous):



OpenStudy (tester97):

\[\frac{ 3 }{ 4 }\times \frac{ 8 }{ 15 }\]  so 2/5 is your answer